OpenStudy (anonymous):

Factor the following trinomial: 15t^2 - 119t + 208

OpenStudy (tkhunny):

If it can be factored, the roots will be symmetric around t = 119/(2*15) = 119/30 = 4 \(119^{2} - 4*15*208 = 41^{2}\) -- Okay, you can factor it. Go!

OpenStudy (anonymous):

(3 t-16) (5 t-13)

OpenStudy (anonymous):

Thank you!